Sol Summary – April 10th

Crew 314 Sol Summary Report 10-04-2025

Sol: 4

Summary Title: Mars the Beautiful

Author’s name: Arnaud de Wergifosse (Crew executive officer)

Mission Status: ongoing

Sol Activity Summary: This morning, Béatrice (commandant), Antoine (engineer), and Arnaud (executive officer) headed out on an EVA. After a quick check of Antoine’s equipment, the crew took Galileo Road toward Candor Chasma, a stunning canyon with striking Martian-like scenery. Dressed in their space suits, they may have felt (just for a few moments) like true explorers on Mars. The EVA went smoothly and offered breathtaking views throughout.

Back at the MDRS, the afternoon was more relaxed, focused on scientific work. Arnaud and Béatrice continued data collection for their joint experiment, which tracks heart rate, heart rate variability, sleep quality and duration, as well as responses to a supplementation protocol (either a placebo or a combination of Lactobacillus helveticus and glycine). Each participant also carries out a daily five-minute cardiac coherence session, a breathing technique known to support stress management.

Meanwhile, our journalist, Batoul continued her research by placing bacterial cultures outside to observe how extreme temperature variations affect their growth, an experiment that may offer insights into the effects of climate change. Bérengère, the greenhab officer, spent the whole day sterilizing spores for her experiment.

To wrap up the day, some crew members took time to unwind or get active. Arnaud finished his afternoon with a high-intensity bike session, a powerful way to cap off a day of exploration, research, and physical challenge.

Look Ahead Plan: One EVA is scheduled in the morning to conduct Bérengère’s experiment. Throughout the day, we will continue working on our individual experiments, including a CRM simulation in the afternoon involving the whole team for Odile’s experiment. Béatrice and Arnaud will carry on their research while collecting data with their wearables.

Anomalies in work : None was mentioned today.

Weather: Sunny, without clouds. Warm in the afternoon. Temperatures varying between 40°and 84° throughout the day

Crew Physical Status : Everyone was in good health today, with no issues reported.

EVA: This morning’s EVA crew changed slightly as Batoul Tani (Crew Journalist) had to remain at the Hab due to a scientific emergency related to her experiment.

The team departed on time, with Arnaud as EVA leader. He drove Curiosity with Antoine (Crew Engineer) as passenger, while Béatrice (Crew Commander) drove Opportunity, since Spirit is still out of use due to a battery charging issue. Their first stop was at the intersection of Galileo Road and Cow Dung Road, after taking Moonwalk Road out of the MDRS. Antoine’s equipment was checked and found to be in good condition.

They continued toward Compass Rock, their intended exploration site. A moment of concern arose due to a miscommunication with Mission Support, who briefly thought the team might be stuck. The situation was quickly clarified, allowing the EVA to proceed.

A brief stop along Galileo Road followed, though no one exited the rover. Communication with CapCom was attempted but hampered by a poor signal. At that point, leadership was transferred to Béatrice. The team’s actual position slightly differed from the initial plan, but they reached Compass Rock by 10:15. After reviewing the map and planned route through the canyon (Candor Chasma), they left the rovers at 10:25.

Due to the lack of signal in Candor Chasma, there was no contact with the team while they were walking. Communication was restored at 10:53 as they exited the canyon.

On the return trip, Antoine drove Curiosity solo, and Arnaud drove Opportunity with Béatrice as passenger. The crew returned safely to the Hab and enjoyed spotting a few lizards along the way.

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, Crew Photos, EVA Report, EVA Request

Support Requested: none

Sol Summary – April 2025

Crew 314 Sol Summary Report 11-04-2025

Sol: 5

Summary Title: Life Between Red Dust and Blue Skies

Author’s name: Arnaud de Wergifosse (Crew executive officer)

Mission Status: ongoing

Sol Activity Summary: Early in the day, an EVA was carried out to collect soil samples for Bérengère’s GreenHab experiment and to check on the equipment for Antoine’s ongoing engineering project. Both tasks were completed successfully, despite the hot conditions outside.

Back in the station, Bérengère planted new seeds in the GreenHab: radish sprouts, arugula, cress, and other fast-growing greens. She also took time to water the existing plants and monitor their development.

Meanwhile, Antoine downloaded the meteorological data, crucial for tracking environmental changes and ensuring the safety of future EVAs.

Batoul, our journalist, spent the afternoon programming a new experiment which simulates the effects of Martian microgravity.

Our daily scientific protocols continue steadily, as each team member follows their assigned experiment schedules with precision and focus.

Life at the MDRS is a constant balance between routine and the unexpected, and today was a perfect example of that.

Look Ahead Plan: One EVA is scheduled in the morning to conduct Antoine’s experiment. Throughout the day, we will continue working on our individual experiments. Béatrice and Arnaud will carry on their research while collecting data with their wearables and with a questionnaire about stress and sleep evaluation. Louis will continue his work in the Observatory.

Anomalies in work: None was notified.

Weather: Sunny without clouds in the morning and the afternoon. Temperatures varied between 44°F and 90°F for the lowest and highest temperatures of the day.

Crew Physical Status: Everyone was in good health today, with no issues reported.

EVA: The EVA team exited the Hab at 09:31 after a 5-minute decompression. Batoul and Odile took Curiosity, while Bérengère and Antoine drove Opportunity. The team passed Cowboy Corner and arrived at the Special Region by 10:05, where they collected two soil samples for Bérengère’s experiment.

Due to poor signal near the region, contact with CapCom was unsuccessful until arrival at the site. Afterward, the team skipped the Overlook due to battery considerations (Opportunity at 58%) and reached the Galileo Road & Cow Dung Road intersection by 11:15. There, they inspected Antoine’s equipment and deployed the drone for additional data.

On the way, they retrieved a fallen rover camera piece. The return began at 11:34 with a driver switch, and the rovers were parked at the Hab by 11:41. Decompression ended at 11:47, marking a successful and efficient EVA.

Reports to be filed: Sol Summary, Journalist Report, Greenhab Report, Operations Report, Crew Photos, EVA Report, EVA Request, Astronomy Report.

Support Requested: Crew requested from the mission support to bring them packages needed for the crew journalist’s experiment

Sol Summary – April 12th

Crew 314 Sol Summary Report 12-04-2025

Sol: 6

Summary Title: Under the Martian Gusts

Author’s name: Arnaud de Wergifosse (Crew Executive Officer)

Mission Status: ongoing

Sol Activity Summary: Despite strong winds and slightly limited visibility, an EVA was successfully conducted today to explore the region. The mission proceeded with caution and allowed for valuable observations of the surrounding terrain.

In the science dome, our crew journalist, Batoul, observed the completion of treatments on her first bacterial cultures. She is expecting her initial results tomorrow, which will mark an exciting milestone in her microbiology experiments. Bérengère, our GreenHab Officer, was able to tend to the plants despite the high temperatures inside the greenhouse, ensuring they remain in good condition.

Overall, it was a relatively calm day for most of the scientific projects, aside from the daily routine experiments which continued as planned.

Look Ahead Plan: One EVA is scheduled in the morning to conduct Bérengère’s experiment at Sea of Shells. Throughout the day, we will continue working on our individual experiments, including a CRM simulation in the afternoon involving the whole team for Odile’s experiment. Béatrice and Arnaud will carry on their research while collecting data with their wearables.

Anomalies in work: None was notified.

Weather: The day was partly cloudy both in the morning and afternoon, with occasional wind and gusts reaching up to 39 mph. Temperatures fluctuated between 47°F and 88°F throughout the entire day.

Crew Physical Status: Everyone was in good health today, with no issues reported.

EVA: Under warmer weather conditions and moderate wind gusts (~15 mph), the crew departed the MDRS at 09:44, heading toward the Sea of Shells for geological sampling and a check on Antoine’s experiment. Despite intermittent signal loss—particularly upon arrival at the Sea of Shells at 10:10—the team successfully completed their tasks. Wind conditions peaked at 16 mph during the outing.

After returning to Brahe Highway at 11:01 and re-establishing communication, the crew made a second stop at the intersection of Galileo and Cow Dung Road to inspect field equipment. The EVA concluded with a safe return to the Hab at 11:24, followed by decompression ending at 11:32.

This EVA successfully met all scientific and operational objectives despite communication and weather-related challenges.

Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA Report, EVA Request, Crew Photos, Astronomy Report.

Support Requested:

A request was submitted to Mission Support to assist the crew in closing the main entrance door, which had suddenly slammed open due to strong winds and became stuck. Thanks to Mission Support’s prompt intervention, the issue was successfully resolved.

Since some of the canned goods and food supplies (such as eggs, coffee, etc.) were already partially used upon our arrival, would it be possible to receive a resupply?

Sol Summary – April 13th

Crew 314 Sol Summary Report 04-13-2025

Sol : 7

Summary Title : Whispers of a forgotten ocean.

Author’s name: Arnaud de Wergifosse (Crew Executive Officer)

Mission Status: ongoing

Sol Activity Summary: Sol after sol, the crew continues to push the boundaries of analog Martian exploration with resilience, focus, and scientific curiosity. The morning began with final EVA preparations and a weather check under clear skies and mild winds.

At 09:37, four crew members exited the airlock after standard decompression procedures to carry out a field mission in the Sea of Shells region. Among them was GreenHab Officer Bérengère Bastogne, who aimed to collect two soil samples from this geologically rich area for her fungi-based research. The site, thought to have once been covered by an ancient Martian ocean, offered both symbolic and scientific value. She was joined by Crew Engineer Antoine, tasked with inspecting his deployed equipment after it had been exposed to intense wind conditions the day prior.The EVA team was split between two rovers: Curiosity (Arnaud and Bérengère) and Opportunity (Antoine and Béatrice). Shortly after departure, Mission Support issued a reminder to reduce rover speed in order to conserve battery for the full route.

After checking in at several intersections, the team reached the vicinity of their target area around 10:20. While on site, they conducted geological sampling, explored the surrounding terrain, and deployed the drone to capture aerial imagery. Due to the local topography, radio contact with the Hab was lost after 09:57, though GPS tracking remained active throughout the EVA. The team began their return journey at 11:20 and successfully re-established communication at 11:40 upon reaching the Cow Dung Road and Galileo Road intersection. There, they performed a brief inspection of Antoine’s equipment and shared a status update with CapCom.

The EVA concluded with the crew safely returning to the MDRS at 11:55. Antoine and Béatrice alternated as drivers for Opportunity, while Arnaud continued as driver of Curiosity. All mission objectives were accomplished, marking another efficient and productive sol of surface operations.

Later in the day, the crew took time to unwind with a group workout session, helping to relieve both the physical fatigue and mental tension accumulated throughout the sol. This was followed by a communal cooking session, where the team shared a warm meal—an enjoyable and grounding end to a demanding but fulfilling day.

Look Ahead Plan : Tomorrow, we will carry out Louis’s experiment, which requires the entire crew to participate in an EVA, meaning there will be two EVAs : one in the morning and another in the afternoon. His experiment, which began on Sol 4 inside the MDRS and will continue in the field on Sol 8, focuses on the interaction between humans and technology. For this purpose, he needs the crew to perform a series of movements, which will be recorded by a wrist-mounted portable device.

Anomalies in work: None was notified.

Weather: The day began with clear, sunny conditions in the morning, followed by increasing cloud cover in the afternoon. Temperatures ranged from 38°F to 79°F over the course of the day.

Crew Physical Status: Everyone was in good health today, with no issues reported.

EVA: Four crew members took part in today’s EVA, including GreenHab Officer Bérengère Bastogne, who collected two soil samples from the Sea of Shells as part of her fungi-related experiment. This site, once covered by an ancient ocean, offered a poetic backdrop to the mission. She was joined by Crew Engineer Antoine, who needed to inspect his installation following strong winds the previous day. The team left the MDRS at 09:37 after standard decompression, splitting into two rovers: Béatrice and Antoine in Opportunity, and Bérengère with EVA leader Arnaud in Curiosity. Mission Support reminded the team early on to reduce speed to preserve rover battery life.

The crew almost reached their destination at 10:20 after checking in at key intersections. While on site, they explored the area, collected samples, and used the drone to capture aerial images. Radio contact was lost after 09:57 due to poor signal but GPS tracking remained active. At 11:20, the team began their return and re-established contact at 11:40, upon arriving at the Cow Dung Road and Galileo Road intersection. There, they checked Antoine’s equipment and briefly reported their activities. The EVA concluded with the team safely heading back, with Antoine and Béatrice alternating driving duties in Opportunity, and Arnaud remaining at the wheel of Curiosity.

Crew was back at the station at 11:55.

Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA Report, EVA Request, Crew Photos, Astronomy Report.

Support Requested: Food requested by the crew was provided by the mission support

Sol Summary – April 14th

Crew 314 Sol Summary Report 14-04-2025

Sol: 8

Summary Title: Sol by Sol, We Explore Further

Author’s name: Arnaud de Wergifosse (Crew executive officer)

Mission Status: ongoing

Sol Activity Summary: Today was marked by scientific advancement and teamwork. Bérengère observed early mycelium development in her fungal experiment, while Batoul obtained initial results from her microbiology study. Two EVAs supported Louis’s astronomy experiment, with Antoine scouting new sites and retrieving equipment. Odile worked on outreach content, and both Bérengère and Batoul conducted microscopic analyses in the Science Dome. The day concluded with a workout and a team meeting to plan upcoming tasks. Another productive day on Mars!

Look Ahead Plan: Tomorrow morning, we will go on EVA to put Antoine’s device back, since we weren’t able to install it during today’s EVA due to the area being unsuitable. We’ll also collect soil samples for Bérengère. Meanwhile, the rest of the crew will continue working on their respective experiments.

Anomalies in work : None was notified.

Weather: The weather was sunny in the morning and the afternoon. Temperatures varied between 39°F and 77°F during the entire day.

Crew Physical Status: Cf HSO report.

EVA:

Two EVAs were successfully conducted today under optimal weather and operational conditions. Both EVAs achieved their primary objectives and adhered strictly to safety and procedural standards.

EVA 1 #10 : The team began decompression at 9:32 AM and exited the habitat at 9:37 AM. Using Spirit and Curiosity, they arrived at the designated site by 9:47 AM. The primary objective was to collect individual movement data for an ongoing experiment on human motion in EVA suits. The team departed the site at 10:42 AM and re-entered the station by 10:55 AM. All procedures were followed without issue.

EVA #11 : The second EVA began at 2:09 PM. The crew used Opportunity and Perseverance, reaching the first stop at Galileo Road to conduct an experiment on human-technology interaction. Afterward, they proceeded to Cowboy Corner to scout a new location for a technical installation. Due to unsuitable soil conditions, the equipment was brought back to the station. The EVA concluded with airlock re-entry at 3:41 PM. Operations were smooth and safely executed.

Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA Report, EVA Request, Crew Photos, and, if applicable, the Astronomy Report.

Support Requested: Crew was wondering if Mission Support would be interested in having a stationary bike at the station for future crews, particularly for potential physical activity research. The bike can be disassembled and stored in its original box, along with the necessary tools for easy reassembly.

Sol Summary – April 15th

Crew 314 Sol Summary Report 15-04-2025

Sol: 9

Summary Title: Sound the Alarm: Emergency Simulation Protocols on Mars

Author’s name: Arnaud de Wergifosse (Executive Officer)

Mission Status: Ongoing

Sol Activity Summary: Today’s mission blended scientific work, routine duties, and a high-stress emergency drill. The day began with a successful EVA by Bérengère, the GreenHab Officer, who collected soil samples under favorable conditions. That night, the team underwent a surprise medical emergency simulation, testing their coordination and stress response in challenging conditions. Bérengère also managed greenhouse tasks and handled the team’s finances. The day’s events were captured in photos showing moments of action, fatigue, and camaraderie.

Look Ahead Plan: Tomorrow, two EVAs are scheduled. The first will take place in the morning, during which Bérengère will collect soil samples from both Candor Chasma and Green Mars View, as the samples gathered today were of poor quality. The second EVA is planned for the afternoon and will involve exploring the area around Skyline Rim to take photographs and document the terrain. The crew will also continue to work on their experiment.

Anomalies in work: None was notified today.

Weather: Partly cloudy in the morning and in the afternoon. Temperatures varied between 42°F and 77°F during the entire day

Crew Physical Status: Everyone was in good health today, with no issues reported.

EVA: The EVA team began decompression at 9:34 AM and exited the Hab by 9:39 AM. After a short walk north of the Science Dome, they identified a new site for Antoine’s experiment and completed equipment setup by 10:20 AM. The crew then resumed their journey using Curiosity (driven by Antoine with Odile) and Perseverance (driven by Béa with Bérengère). A brief communication stop was made at 10:46 AM before reaching Green View. The return trip began at 11:48 AM with a switch in drivers. The team safely arrived at the MDRS at 12:12 PM, and decompression was completed by 12:20 PM. The EVA was successfully conducted without incident.

Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA Report, EVA Request, Astronomy Report.

Support Requested: /

Sol Summary – April 16th

Crew 314 Sol Summary Report 16-04-2025

Sol: 10

Summary Title: Martian Simulation, Human Reality.

Author’s name: Arnaud de Wergifosse (Crew executive officer)

Mission Status : ongoing

Sol Activity Summary: Today, two EVAs were conducted.

The first EVA took place in Condor Chasma and was primarily aimed at collecting soil samples for Bérengère (GreenHab Officer) as part of her ongoing microbiological experiment. The mission went smoothly under favorable weather conditions.

The second EVA, conducted at Skyline Rim, proved to be more challenging due to strong winds which significantly hindered movement and visibility. Despite the adverse conditions, the crew managed to achieve the primary objective: exploring and assessing new terrain that could be suitable for future EVAs.

Meanwhile, in the Science Dome, Batoul (Journalist) analyzed the results of the Bacillus thuringiensis experiment. The initial outcomes appear promising and will undergo further analysis for interpretation. In addition, several other experiments reached completion today, with samples either archived or disposed of in accordance with the established protocols.

In the main habitat, the end of the mission is beginning to be felt. Louis (Astronomer) has started analyzing the data from his observations, and several crew members have begun working on their End of Mission Reports.

In the lab, Bérengère carried out a staining procedure on fungal samples using violet dye to assess germination. The dye improved visibility of the spores under the microscope, and preliminary observations indicate successful germination in some samples.

Despite the difficult weather conditions encountered during the second EVA, the day was productive both in the field and in the laboratory.

Look Ahead Plan : One final EVA is scheduled for the morning to both take down Antoine’s experiment, due to strong winds and rain expected from tomorrow afternoon through Friday, and to visit Green Mars View. With only two days left at the station, the rest of the crew will focus on wrapping up and finalizing their experiments.

Anomalies in work: None was notified.

Weather: Partly cloudy during the morning and the afternoon. Temperatures varied between 51°F and 80°F during the entire day

Crew Physical Status: Everyone was in good health today, with no issues reported.

EVA: There were two EVA today.

The first EVA began at 9:34 AM with the standard five-minute decompression. The team departed the MDRS at 9:43 AM in two rovers—Spirit (driven by Bérengère with Odile) and Opportunity (driven by Batoul with Louis). After briefly overshooting their target by 500 meters, they corrected course and arrived at the designated site at 10:08 AM. The crew proceeded on foot into the canyon at 10:11 AM, reaching it by 10:23 AM. A short communication loss occurred around 10:54 AM. The team began their return at 11:02 AM, reaching the rovers by 11:28 AM and arriving back at the MDRS at 11:42 AM. The EVA concluded with decompression ending at 11:49 AM. The mission was completed successfully.

The second EVA began decompression at 1:57 PM and exited the station at 2:03 PM, heading north around the Science Dome to inspect Antoine’s field equipment. They reached the site at 2:05 PM and departed two minutes later, continuing toward Skyline. Although communication was disrupted due to poor signal, contact was restored at 2:53 PM.

Weather conditions included sustained winds of 8 mph with gusts up to 18 mph. As conditions worsened around 3:35 PM, the team decided to end the EVA. The terrain, while mostly flat, poses access challenges due to undefined paths and elevation changes, making rover use difficult.

The EVA concluded with repressurization from 4:12 to 4:17 PM. Despite the wind, the area remains suitable for future EVAs in better weather.

Reports to be filed: Sol Summary, Journalist Report, Crew Photos, GreenHab Report, Operations Report, EVA Report, EVA Request, Astronomy Report.

Support Requested: None

Sol Summary – April 17th

Crew 314 Sol Summary Report 04-17-2025

Sol: 11

Summary Title: As the Mission Winds Down

Author’s name: Arnaud de Wergifosse

Mission Status: Ongoing

Sol Activity Summary: Despite extreme weather conditions, including rain and lightning, the crew managed to complete a morning EVA to retrieve Antoine’s equipment. All other outdoor activities were canceled, including tomorrow’s EVA. The crew used the downtime to focus on filming and finalizing scientific work. Batoul completed lab work for her microbiology experiment, and the final emergency simulation—a smoke-filled Greenhab drill—was led by Odile. Physical activity remained intense as Béa and Arnaud’s cognitive performance study nears completion. Bérengère continued advancing her Greenhab experiment as the mission enters its final days.

Look Ahead Plan: No EVAs are planned for tomorrow due to the weather conditions. The ground is expected to be wet, making it difficult for both rover travel and walking. Instead, we’ll focus on finalizing our experiments and beginning to pack up, as it will already be our last Sol at the station.

Anomalies in work: None was notified.

Weather: Sunny with a few clouds for the morning and the afternoon. Windy after 11am with wind gusts attaining 37mph and a constant wind between 16 and 21 mph. Temperatures varied between 35°F and 71°F during the entire day

Crew Physical Status: Everyone was in good health today, with no issues reported

EVA: The EVA team departed the MDRS at 9:11 AM, with Béatrice driving Curiosity (Antoine as passenger) and Louis driving Perseverance (Arnaud as passenger). After a brief stop in the Special Region at 9:25 AM, the crew arrived at Green Mars View at 9:48 AM, where they collected soil samples for Bérengère’s experiment.

At 10:20 AM, the team left the site with Arnaud now driving Perseverance and Antoine taking over Curiosity. They returned to the MDRS at 10:45 AM and reported rover battery usage to CapCom. At 10:49 AM, the crew proceeded on foot to dismantle Antoine’s experiment and bring the equipment back to the station.

Reports to be filed: Sol Summary, Journalist Report, Crew Photos, GreenHab Report, Operations Report, EVA Report, EVA Request, Astronomy Report (if applicable).

Support Requested: None

Sol Summary – April 18th

Crew 314 Sol Summary Report 18-04-2025

Sol: 12

Summary Title: The Final Sol: Closing a Chapter on Simulated Mars

Author’s name: Arnaud de Wergifosse (Crew executive officer)

Mission Status: Ending (04-19-2025 at 6 AM).

Sol Activity Summary: On the final day of the mission, the crew embraced their remaining tasks with a bittersweet mix of motivation and nostalgia. The day was spent organizing equipment, wrapping up experiments, and capturing final video footage—including a lighthearted segment on cooking with freeze-dried food that reflected their creativity and resilience. Though no EVA could take place due to poor weather, the team maintained their commitment to safety and routine, finishing strong with one last workout. In the Science Dome, Bérengère devoted the day to her research on spore germination and plant-fungus interactions under Martian conditions. As the simulation came to a close, the crew left with a sense of accomplishment, meaningful memories, and a trove of scientific data to bring back to Earth.

Look Ahead Plan: Cleaning, tidying up. Simulation ending tomorrow at 6 AM.

Anomalies in work: None was notified.

Weather: Sunny in the morning and cloudy in the afternoon. Temperatures varied during the entire day between 29°F and 52°F.

Crew Physical Status: Everyone was in good health today, with no issues reported

EVA: No EVA was programed today due to weather conditions

Reports to be filed: Journalist Report and Crew Photos, GreenHab Report, Operations Report, Astronomy Report, Mission summary report

Support Requested: None

Sol Summary – April 21st

Crew 315 Sol Summary Report 21-04-2025

Sol: 1

Summary title: First Day on Mars

Author: Tim Gagnon

Mission status: ongoing

Sol activity summary:

Our day welcomed the rising Sun over the Mars landscape.

As we gathered for breakfast we were saddened to learn that Pope Francis had died.

We enjoyed pancakes prepared by the Crew Artist and discussed our upcoming EVAs.

Anomalies in work: Main generator still does not work, suit 7 doesn’t charge, the greenhab scale doesn’t work

Weather: Sunny with highs in the mid-70s

Crew physical status: We are in good health

EVA: The Engineer, Journalist and HSO conducted the first EVA and we all gathered on the first floor to assist with

suit-up and comm checks.

As this first team completed the pre-breathe our Commander sent them off with him playing the theme from “2001 A Space Odyssey.”

They went to the Marble Ritual and explored that area.

When they returned at 1400 hrs,we conducted an informal debrief and the Commander and Artist prepared for the second EVA.

It was not as smooth and seamless as EVA1 because the Artist struggled with comms and getting into the Curiosity Rover. They traveled to the Marble offering and had less time to explore.

The Commander graciously took photos of the Artist as gifts for friends. The EVA concluded on time at 1600 hrs.

As a child of Apollo during my EVA I was able to imagine what it was like for childhood heroes to walk on the Moon. I even photographed my boot print as Buzz Aldrin did on Apollo 11. It was a glorious experience which I will turn into a fine art post mission.

After the EVAs were completed we met upstairs for a 100cameras photography module conducted by our HSO. The goal of which was to give us a better understanding about composition and help us tell a clearer story with our photography. The class was made up of instruction and then a photo exploration of our campus illustrating key words.

Our HSO and journalist prepared a delicious dinner of potato and vegetable soup and homemade bread. Using freeze dried potatoes, broccoli, corn, cheese and herbs harvested from the Green Hab they created a delicious chowder. The homemade bread was made with white and wheat flour and yeast. A bread making machine will be a highly desired appliance on future Mars missions.

After dinner while waiting for the comms window to open, we enjoyed fellowship and spoke of past experiences that we bring to this analog mission. We also consulted the map and discussed possible future EVAs farther away from the Hab.

At this time we are completing our reports, editing photos and preparing to end our day. It was full, it was exciting and we look forward to a good night’s sleep and an interesting day tomorrow.

Look ahead plan: Two EVAs are planned for tomorrow to conduct Michael’s 3D imaging project and other experiments.

Reports to be filed: Sol (filed), journalist report, greenhab report, operations report, crew photos, mission plan (filed)

Support requested: None

Copyright © The Mars Society. All rights reserved. | Main Site